Can I please get some help?
One card is selected at random from a standard 52-card deck of playing cards. Find the probability that the card selected is a red king. (Give solution as a fraction in lowest terms.)


Answer by stanbon(75887)

One card is selected at random from a standard 52-card deck of playing cards. Find the probability that the card selected is a red king.
-----------
There are two red kings in the deck.
P(red king) = 2/52 = 1/26
===========================
Cheers,
Stan H.